OpenStudy (anonymous):

Factor: 36b^2 +60bf +25f^2 = (? ?) (? ?)

OpenStudy (dumbcow):

wow ok find factors of 36*25 that add up to 60

OpenStudy (anonymous):

right, and I feel like I gone through every possible combo and nothing is adding up to 60?!

OpenStudy (anonymous):

25f^2+60bf+36b^2=(5f+6b)(5f+6b) Do you want to know how i got the answer?

OpenStudy (anonymous):

for sure please!

OpenStudy (anonymous):

how the he*& did I miss that one??? ugh. you are fast...

OpenStudy (anonymous):

Well what I did first was rearrange the equation so that f^2 was our leading term from there I used a technique to find my terms in order to get 25f^2 I would have to multiply 5f by 5f

OpenStudy (dumbcow):

oh because the first and last coefficient are perfect squares... it comes out to be a perfect square...

OpenStudy (anonymous):

Now some would figure how to get the second term, but that would actually be a much more tedious process so I moved to 36b^2 So...in order to get 36b^2, i would have to multiply 6b by 6b Now all I have to do combine my terms into binomials to get (5f+6b) And since we are multiplying 5f by 5f and 6b by 6b to get our first and last term, I came up with the binomials (5f+6b)(5f+6b) And if you notice, the product ends up to be... (5f+6b)(5f+6b)=25f^2+30bf+30bf+36b^2 =25f^2+60bf+36b^2

OpenStudy (anonymous):

and you knew to use the factors you used rather than other factors because the terms are squared? or what made you so sure not to use 12*3 for 36, for example?

OpenStudy (anonymous):

how did you avoid the trial and error method is what I am wondering?

OpenStudy (anonymous):

Yeah, I basically found my terms because they were perfect squares. If they weren't I wouldve found them with a different method.